OpenStudy (anonymous):

Kim had a 4% raise. He current salary is $30,420. What was her previous salary?

OpenStudy (anonymous):

i would assign x as the salary before and do: x-0.4x=30,420 0.6x=30,420 and then divide and get the slary before

OpenStudy (amistre64):

30420 = P + P(.004)

OpenStudy (anonymous):

sorry thought it was 40% should be .04

OpenStudy (amistre64):

.04.... :)

OpenStudy (amistre64):

30420/(1.04) = Paycheck last time right?

OpenStudy (amistre64):

29250

OpenStudy (amistre64):

x - .04x = 30420 ??? you sure?

OpenStudy (anonymous):

i meant addition

OpenStudy (amistre64):

ahhh.... i thought you did, I could see it in your eyes.. :)